# Coherent Information Phase Transition in a Noisy Quantum Circuit

### v1

@@ Abstract
Figure 1: Circuit structure and phase diagram. (a) Circuit structure. Orange, green, yellow, and blue rectangles represent unitary gates, measurements, noise, and QE operations, respectively. The initial state is entangled with reference qubits $R$. We primarily consider QE operations where $U_{SA}=\text{SWAP}$. (b) Phase diagram. We focus on the first-order transition from recoverable to irrecoverable phase, tuned by the relative frequency of noise and QE operations.
@@
This statistical mechanics model is a ferromagnetically coupled spin model with random fields oriented in two different directions on a honeycomb lattice [52]. The vertical bond contributes a weight of $\text{Wg}^{Q}(\sigma^{-1}\tau)$, where $\text{Wg}^{Q}(g)$ is the Weingarten function, acting as a ferromagnetic coupling that assigns greater weights to configurations where $\sigma=\tau$. The non-vertical bonds are determined by the specific choice of QE operations and noise. For instance, if we select $U_{SA}=\text{SWAP}$ and consider depolarizing noise, the bond contributes a weight:
$W^{\text{depo}}$ $(\sigma,\tau)=(1-p)(1-q_{t})\left\langle\sigma|\tau\right\rangle+dp(1-qq_{t})+qq_{t}\times$
$[(1-p)\left\langle\sigma|\mathbb{I}\right\rangle+dp]\left\langle\mathbb{I}|\tau\right\rangle+(1-p)(1-q)q_{t}\left\langle\sigma|\mathbb{C}\right\rangle.$
Here, $\left\langle\sigma|\tau\right\rangle=d^{|\sigma\tau^{-1}|}$, where $d$ is the local Hilbert space dimension and where $|\sigma\tau^{-1}|$ denotes the number of cycles in permutation $\sigma\tau^{-1}$. $\mathbb{I}$ corresponds to the identity element in $\mathbb{S}(Q)$. Thus, $q$ determines the net field direction. For small $q$, the net field aligns with $\mathbb{C}$, whereas for large $q$, it aligns with $\mathbb{I}$. When combined with the results in Ref. [52], the full phase diagram is depicted in Fig. 1(b). By identifying $p$ as the temperature and considering noise and QE operations as two competing external fields, the phase diagram is reminiscent of that of a simple 2D Ising model. Consequently, a first-order phase transition is expected when tuning $q$.
@@
Another interesting phenomenon we observed is critical slowing down [63, 64, 65, 66]. As illustrated in Fig. 3(a), the convergence of the coherent information slows significantly near $q=0.5$ for resetting noise. By defining the convergence time $t_{c}$ as the time when $\left|I_{C}(t)-I_{C}(t_{\text{final}})\right|$ first falls below $0.05$, we plot $t_{c}$ as a function of $q$ for different system sizes in Fig. 3(b). The results indicate that $t_{c}$ diverges with system size as $q$ approaches the critical point, suggesting an infinite convergence time in the thermodynamic limit. In contrast, in the absence of unitary gates, while the coherent information still transitions from positive to negative precisely at $q_{c}=0.5$ (corresponding to the scenario where half of the system qubits are discarded), critical slowing down was not observed, as shown in Fig. 3(c) and 3(d). This observation underscores the indispensability of unitary gates for the manifestation of a phase transition.
@@
Efficient protocol. Measuring coherent information for a specific circuit and trajectory requires repeated preparation of the same final state. However, in realistic experiments, noise locations are usually uncontrollable, making it difficult to reliably reproduce the same circuit. This issue is compounded by the post-selection problem when measurements are involved, due to Born’s rule [67, 68, 54]. Motivated by previous work on constructing post-selection-free probes that reflect the correlation between a quantum device and classical simulations [69, 53, 70], we propose a resource-efficient protocol to observe the coherent information phase transition. In the quantum device, an initial state $\rho$ is prepared and encoded by the application of random unitary gates alone. This encoded state then undergoes the circuit, with ancillary qubits introduced during the evolution. The unnormalized final state on $A$ is denoted as $\widetilde{\rho_{A}^{m}}$, with $\mathrm{Tr}(\widetilde{\rho_{A}^{m}})=p_{\rho}^{m}$, where $p_{\rho}^{m}$ represents the probability of the particular trajectory $m$. For brevity, we omit the explicit dependence on the circuit realization $\mathcal{C}$, which includes choices of random unitary gates and locations of noise, QE operations and measurements. Leveraging the knowledge of the measurement outcomes and the circuit realization, we execute a classical simulation of the same circuit, but with the input state being the encoded state of a distinct state $\sigma$. Assuming the unitary gates are Clifford gates and $\sigma$ is a stabilizer state, the classical simulation becomes computationally efficient. We then also obtain $\widetilde{\sigma_{A}^{m}}$, with $\mathrm{Tr}(\widetilde{\sigma_{A}^{m}})=p_{\sigma}^{m}$. It is worth noting that the initial state $\rho$ can be chosen beyond stabilizer states, making the evolution intractable on a classical computer.
@@
Figure 4: Efficient protocol. (a) Schematic workflow of the protocol. (b) Take depolarizing noise as an example. We choose $\rho=(\left|+\right\rangle\left\langle+\right\rangle)^{\otimes L/2}\otimes(\left|0\right\rangle\left\langle 0\right\rangle)^{\otimes L/2}$ and $\sigma=(\left|0\right\rangle\left\langle 0\right\rangle)^{\otimes L}$. Every data point is averaged over $3\times 10^{3}$ realizations. The inset shows the data collapse result. (c) Analogous to (b), with measurement probability $p=0.1$.
@@
Physically, this probe reflects the distinguishability of different initial states by leveraging access to both the measurement outcomes and the final state of the ancilla qubits. In the recoverable phase, the initial information is preserved, and we therefore expect $\chi=0$, corresponding to perfect distinguishability. Conversely, in the irrecoverable phase, the initial information becomes erased, causing $\chi$ to approach 1. A detailed analytical analysis of this protocol is provided in [62]. To demonstrate the efficacy of this probe, we present numerical results in Fig. 4(b) and Fig. 4(c) for the case of depolarizing noise with $p=0$ and $p=0.1$, respectively. The critical points are determined to be $q_{c}=0.362(2)$ and $q_{c}=0.378(2)$, which are in agreement with the exact phase transition points.
@@ Acknowledgements.
Acknowledgment. We are deeply indebted to Xiao-Liang Qi for many valuable discussions. This work is supported by the National Key Research Program of China under Grant No. 2019YFA0308404, the Natural Science Foundation of China through Grants No. 12350404 and No. 12174066, the Innovation Program for Quantum Science and Technology through Grant No. 2021ZD0302600, the Science and Technology Commission of Shanghai Municipality under Grants No. 23JC1400600 and No. 2019SHZDZX01.

### v2

@@ Abstract
Figure 1: Circuit structure and phase diagram. (a) Circuit structure. Orange, green, yellow, and blue rectangles represent unitary gates, measurements, noise, and QE operations, respectively. The initial state is entangled with reference qubits $R$. We primarily consider QE operations where $U_{SA}=\text{SWAP}$. (b) Phase diagram. We focus on the first-order transition from recoverable to irrecoverable phase, tuned by the relative frequency of noise and QE operations. This phase transition manifests irrespective of the measurement rate $p$.
@@
This statistical mechanics model is a ferromagnetically coupled spin model with random fields oriented in two different directions on a honeycomb lattice [52]. The vertical bond contributes a weight of $\text{Wg}^{Q}(\sigma^{-1}\tau)$, where $\text{Wg}^{Q}(g)$ is the Weingarten function, acting as a ferromagnetic coupling that assigns greater weights to configurations where $\sigma=\tau$. The non-vertical bonds are determined by measurements, QE operations and noise. While their exact contributions depend on the specific QE operations and noise model, QE operations generally behave as external fields applied at random locations, aligning in the direction $\mathbb{I}$, corresponding to the identity element in $\mathbb{S}(Q)$. Conversely, noise acts as an external field oriented in the direction $\mathbb{C}$. Concrete expressions for the bond weights under various noises are provided in Ref. [59]. The parameter $q$ governs the relative frequency of QE operations and noise, thereby tuning the net field direction. For small $q$, the net field aligns with $\mathbb{C}$, whereas for large $q$, it aligns with $\mathbb{I}$. When combined with the results in Ref. [52], the full phase diagram is depicted in Fig. 1(b). By identifying $p$ as the temperature and considering noise and QE operations as two competing external fields, the phase diagram is reminiscent of that of a simple 2D Ising model. Consequently, a first-order phase transition is expected when tuning $q$.
@@
Another interesting phenomenon we observed is critical slowing down [63, 64, 65, 66]. As illustrated in Fig. 3(a), the convergence of the coherent information slows significantly near $q=0.5$ for resetting noise. By defining the convergence time $t_{c}$ as the time when $\left|I_{C}(t)-I_{C}(t_{\text{final}})\right|$ first falls below $0.05$, we plot $t_{c}$ as a function of $q$ for different system sizes in Fig. 3(b). The results indicate that $t_{c}$ diverges with system size as $q$ approaches the critical point, suggesting an infinite convergence time in the thermodynamic limit. In contrast, in the absence of unitary gates, while the coherent information still transitions from positive to negative precisely at $q_{c}=0.5$ (corresponding to the scenario where half of the system qubits are discarded), critical slowing down was not observed, as shown in Fig. 3(c) and Fig. 3(d). This observation underscores the indispensability of unitary gates for the manifestation of a phase transition.
@@
Efficient protocol. Measuring coherent information for a specific circuit and trajectory requires repeated preparation of the same final state. However, in realistic experiments, noise locations are usually uncontrollable, making it difficult to reliably reproduce the same circuit. This issue is compounded by the post-selection problem when measurements are involved, due to Born’s rule [67, 68, 54]. Motivated by previous work on constructing post-selection-free probes that reflect the correlation between a quantum device and classical simulations [69, 53, 70, 71, 72], we propose a resource-efficient protocol to observe the coherent information phase transition. In the quantum device, an initial state $\rho$ is prepared and encoded using random unitary gates alone. This encoded state then undergoes the circuit, incorporating ancilla qubits during the process. The unnormalized final state on $A$ is denoted as $\widetilde{\rho_{A}^{m}}$, with $\mathrm{Tr}(\widetilde{\rho_{A}^{m}})=p_{\rho}^{m}$, where $p_{\rho}^{m}$ represents the probability of the particular trajectory $m$. For brevity, we omit the explicit dependence on the circuit realization $\mathcal{C}$, which includes choices of random unitary gates and locations of noise, QE operations and measurements. Leveraging the knowledge of the measurement outcomes and the circuit realization, we simulate the circuit classically with an encoded version of a distinct input state $\sigma$. Assuming the unitary gates are Clifford gates and $\sigma$ is a stabilizer state, the classical simulation becomes computationally efficient. We then also obtain $\widetilde{\sigma_{A}^{m}}$, with $\mathrm{Tr}(\widetilde{\sigma_{A}^{m}})=p_{\sigma}^{m}$ [73].
@@
Figure 4: Efficient protocol. (a) Schematic workflow of the protocol. An important additional ingredient, compared to the approach in [53], is the incorporation of projective measurements on the ancilla qubits. (b) Take depolarizing noise as an example. We choose $\rho=(\left|+\right\rangle\left\langle+\right\rangle)^{\otimes L/2}\otimes(\left|0\right\rangle\left\langle 0\right\rangle)^{\otimes L/2}$ and $\sigma=(\left|0\right\rangle\left\langle 0\right\rangle)^{\otimes L}$. Every data point is averaged over $3\times 10^{3}$ realizations. The inset shows the data collapse result. (c) Analogous to (b), with measurement probability $p=0.1$.
@@
The proxy $\chi$ resembles the “quantum” cross entropy introduced in Ref. [53], with the notable distinction that ancilla qubits are introduced dynamically in this context. Furthermore, while this proxy does not reveal MIPT, it proves to be a robust indicator of the coherent information transition. As demonstrated in Supplemental Material [59], $\chi$, akin to the coherent information, is directly related to the free energy difference between distinct boundary conditions imposed on the bottom boundary. This dependency drives a sharp transition in $\chi$ from 00 to $1$. Physically, this probe reflects the distinguishability of different initial states by leveraging access to both the measurement outcomes and the final state of the ancilla qubits. In the recoverable phase, the initial information is preserved, and we therefore expect $\chi=0$, corresponding to perfect distinguishability. Conversely, in the irrecoverable phase, the initial information becomes erased, causing $\chi$ to approach 1.
To demonstrate the efficacy of this probe, we present numerical results in Fig. 4(b) and Fig. 4(c) for the case of depolarizing noise with $p=0$ and $p=0.1$, respectively. The critical points are determined to be $q_{c}=0.362(2)$ and $q_{c}=0.378(2)$, which are in agreement with the exact phase transition points.
@@ Acknowledgements.
Acknowledgment. We are deeply indebted to Xiao-Liang Qi for many valuable discussions. This work is supported by the Natural Science Foundation of China through Grants No. 12350404 and No. 12174066, the Innovation Program for Quantum Science and Technology through Grant No. 2021ZD0302600, the Science and Technology Commission of Shanghai Municipality under Grants No. 23JC1400600, No. 24LZ1400100 and No. 2019SHZDZX01.